OpenStudy (anonymous):

Two baskets contains snacks. The baskets contain 15 bags of peanuts, of which 6 are reduced sodium and 8 granola bars, of which 2 are reduced sodium. If you randomly select one snack from each basket, what is the probability that both snacks are reduced sodium and 1 snack is a bag of peanuts and the other is a granola bar?

OpenStudy (anonymous):

p = p(low sodium nuts then low sodium granola) + p(low sodium granola then low sodium nuts) = 6/23 x 2/22 + 2/23 x 6/22 = 24/506 = 12/253